OpenStudy (anonymous):

Find the point P on the graph of the function y=sqrt{x} closest to the point (3,0)

OpenStudy (anonymous):

'The x coordinate of P is:'

OpenStudy (ranga):

Find the distance of the point (x,y) on the curve to the point (3,0) substitute y with sqrt(x) Minimize the distance by finding the derivative with respect to x, equating it to 0 and solving for x.

OpenStudy (ranga):

Distance between (x,y) and (3,0) is: D = sqrt{ (x-3)^2 + y^2 } y = sqrt(x) y^2 = x put this in D D = sqrt{ (x-3)^2 + x } = sqrt(x^2 - 6x + 9 + x} D = sqrt(x^2 - 5x + 9) Minimize D by finding the derivative with respect to x, equating it to 0 and solving for x.

OpenStudy (ranga):

answering someone else. will be right back.

OpenStudy (anonymous):

Thank you

OpenStudy (ranga):

D = sqrt(x^2 - 5x + 9) To minimize, find derivative and set it to 0. D' = 1/2(x^2 - 5x + 9)^(-1/2) * (2x - 5) = 0 D' will be zero when numerator = 0 2x - 5 = 0 x = 5/2 Put x = 5/2 back in y=sqrt{x} y = sqrt(5/2) So the point on the curve that is closest to the point (3,0) is: (5/2, sqrt(5/2))

OpenStudy (ranga):

In decimal, the point will be (2.5, 1.58)

OpenStudy (anonymous):

Thank you so much, I'm going to "fan" you haha

OpenStudy (ranga):

you are welcome. glad to be able to help.